OpenStudy (anonymous):

how do you get 9x=12-6y into slope-intercept form?

OpenStudy (azureilai):

When trying to get slope intercept form, you want to isolate y to one side of the equation. essentially, the formula is y=mx+b So 9x=12-6y -6y+9x=12 -6y=9x+12 Now divide everything by (-6) to simplify

OpenStudy (anonymous):

so y= -1.5+2?

OpenStudy (anonymous):

1.5x*

OpenStudy (azureilai):

You were right the first time. it is y=-1.5x+2 I made a mistake in writing the equation the equation before you divided everything was suppose to be 6y=-9x+12 I mixed up my signs.